# Reassessing How to Compare and Improve the Calibration of Machine Learning Models

### v1

@@ 4.1 General Theory
$\mathbb{E}[d_{\phi}(\mathbb{E}[Z\mid X],X)]$ $=\mathbb{E}[\phi(\mathbb{E}[Z\mid X])]-\mathbb{E}[\phi(X)]+\mathbb{E}[\nabla\phi(X)^{\top}(\mathbb{E}[Z\mid X]-X)]$
$=\mathbb{E}[\phi(\mathbb{E}[Z\mid X])]-\mathbb{E}[\phi(X)]+\mathbb{E}[\nabla\phi(X)^{\top}(Z-X)].$ (4.3)
@@
Although bias-variance decompositions have appeared for Bregman divergences (Pfau, 2013) and calibration-sharpness decompositions have appeared for proper scoring rules (Kull & Flach, 2015; Pohle, 2020), we are not aware of Lemma 4.1 being used in the literature. We note that both the sharpness gap term and the calibration term in (4.3) are nonnegative, by Jensen’s inequality.
Lemma 4.1 has several consequences. Firstly, we obtain a notion of calibration error from any choice of Bregman divergence, since the calibration term in (4.2) is 0 if and only if $E[Z\mid X=x]=x$ a.e. This means when we specify a (Bregman) loss, we are implicitly also specifying a calibration error, and vice versa. We can thus use the overall loss associated with a predictor that achieves zero calibration error to detect “trivial calibration”. To illustrate this, we can show that the MRR strategy from the previous section - which produces a constant calibrated predictor - is in fact the worst possible perfect calibration strategy from the perspective of expected loss.
Theorem 4.2.
The proof of this result requires the following proposition, a version of which also appeared in Pohle (2020) for the decomposition therein.
Proposition 4.3.
For $g_{1},g_{2}:\mathbb{R}^{d}\to\Delta^{k-1}$ and $\pi$ as in Lemma 4.2, if $\sigma(g_{1}(X))\subseteq\sigma(g_{2}(X))$, then:
@@
Practical ramifications. The point of proving Theorem 4.2 is to show that if we have an appropriate Bregman divergence $d_{\phi}$ that is associated with our calibration error, we can compare calibrated models using $d_{\phi}$ to detect differences that metrics that depend only on $c(X)$ will not be sensitive to. Furthermore, Proposition 4.3 shows that the sharpness gap term in Lemma 4.1 captures the intuitive idea that the more “fine-grained” a calibrated model is, the better its generalization.
As mentioned earlier, the most reported calibration errors are versions of the ECE. It is easy to see from Lemma 4.1 that the squared $L^{2}$ ECE is the calibration error induced by MSE/Brier score, i.e. $d_{\phi}(x,y)=\norm{x-y}^{2}$. There is however, no straightforward correspondence for the $L^{1}$ ECE (since absolute value is not a Bregman divergence). Based on this we can make the following recommendation: if choosing to report ECE, it makes more sense to report the $L^{2}$ ECE alongside MSE/Brier score (or just the leftover sharpness gap).
4.2 Adapting to Confidence Calibration, and Additional Pitfalls
The above discussion relies on the full calibration notion of (2.1), as the calibration term in Lemma 4.1 introduces $\mathbb{E}[Y\mid g(X)]$. However, as we have mentioned, for multi-class classification we rely on the notion of confidence calibration for efficiency reasons. Thus, it is natural to ask how to best adapt the previous observations to the confidence calibration setting.
One approach is to apply a 1-D version of our choice of $d_{\phi}$ to the binary classification problem derived from applying the standard confidence calibration transformation, so that all of the previous results directly translate. More precisely, we replace the multi-class label $Y$ with the binary label $\mathbbm{1}_{Y=c(X)}$ and we replace the probability vector $g(X)$ with the scalar confidence $h(X)$. In practice, this is estimated by computing the mean of $d_{\phi}$ on (accuracy, predicted confidence) tuples obtained from test data.
This approach appears in the literature; for example, in the very recent work of Tian et al. (2023) surveying strategies for eliciting confidence scores from LLMs. Unfortunately, we point this out as another pitfall - the loss $d_{\phi}$ incurred on the confidence calibration problem can be arbitrarily small for a predictor that is always wrong despite such a predictor having true loss on the multi-class classification problem bounded below by a constant. The next proposition illustrates this for MSE.
Proposition 4.4.
The proof follows from directly evaluating (4.6). Fortunately, in almost all cases of interest, estimating $\mathbb{E}[d_{\phi}(\pi_{Y\mid X},g(X))]$ is computationally efficient. We thus advocate for reporting the divergence in the multi-class setting and then the calibration error in the confidence calibration setting, which is also a relatively standard practice. This is further justified in that, under reasonable conditions on $d_{\phi}$ and its 1-D counterpart, the confidence calibration error is a lower bound on the full calibration error in the decomposition from Lemma 4.1.
Lemma 4.5.
For any Bregman divergences $d_{\phi}:\Delta^{k-1}\to\overline{R}$ and $d_{\varphi}:[0,1]\to\overline{\mathbb{R}}$ satisfying $d_{\phi}(x,y)\geq d_{\varphi}(x_{i},y_{i})$ for all $i\in[k]$, it follows that:
$\mathbb{E}[d_{\phi}(\mathbb{E}[\pi_{Y\mid X}\mid g(X)],g(X))]\geq\mathbb{E}[d_{\varphi}(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)],h(X))].$ (4.7)
Corollary 4.6.
The inequality (4.7) still holds under the conditions that $d_{\phi}$ satisfies the data-processing inequality and $d_{\varphi}$ is obtained by applying a transition kernel to the arguments of $d_{\phi}$.
The proof of the above relies on similar techniques to that of Lemma 4.1, i.e. a careful analysis of the conditional expectation terms. Lemma 4.5 covers MSE/Brier score since we can take $\phi(x)=\norm{x}^{2}$ and $\varphi(x)=x^{2}$, while Corollary 4.6 covers KL since the logistic loss $\varphi(x)=x\log(x)+(1-x)\log(x)$ can be obtained from applying a transition kernel depending only on a chosen coordinate of $x$.
5 Calibration-Sharpness Diagrams
To summarize our observations thus far, we have recommended that calibration should be compared using the calibration metric associated with the Bregman divergence for which the models were optimized, and that both should be reported together to avoid trivial cases of calibration. It remains to address how to similarly modify the visualization of (confidence) calibration, which is typically done via reliability diagrams that plot $\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)]$ for different values of $h(X)$. Our goals are two-fold: we want a visualization that is efficient to compute, and also makes it obvious whether one model is better calibrated and/or has better overall divergence than another model.
The approach we take is to augment the plot of $\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)]$ with a surrounding band whose length $\rho(h(X))$ corresponds to the sharpness gap from Lemma 4.1 at $h(X)$, which we refer to as calibration-sharpness diagrams. Now letting $d_{\phi}$ denote our Bregman divergence and $d_{\varphi}$ denote its appropriately defined 1-D counterpart as in the previous section,
$\rho(h(X))=\mathbb{E}\left[d_{\phi}(\pi_{Y\mid X},g(X))\mid h(X)\right]-d_{\varphi}\left(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)],h(X)\right).$ (5.1)
In the binary classification case, (5.1) exactly corresponds to a pointwise version of the sharpness gap term from Lemma 4.1. In the multi-class setting, it no longer has the same interpretation since we are subtracting the confidence calibration error as opposed to the full calibration error. However, we are guaranteed that the band length is nonnegative (at least for standard choices of $\phi$) thanks to Lemma 4.5, and in this setting the band can be interpreted as capturing both the sharpness gap as well as the difference between full and confidence calibration. The benefit of the formulation in (5.1) is that both terms require estimating only 1-D conditional expectations, which can be done efficiently.
For our estimation procedure, we use Nadaraya-Watson kernel regression (Nadaraya, 1964; Watson, 1964) with a kernel $K$ and bandwidth hyperparameter $\sigma$ following the estimation approaches of Błasiok & Nakkiran (2023) and Chidambaram et al. (2024). We recall that given $n$ data points $(x_{i},y_{i})\in\mathbb{R}\times\mathbb{R}$ the kernel regression estimator of $\mathbb{E}[Y\mid X=x]$ is defined as:
Previous work (Zhang et al., 2020; Popordanoska et al., 2022; Chidambaram et al., 2024) has shown that we can obtain consistent estimators of $L^{p}$ ECE using (5.2). We extend these results to calibration errors derived from Bregman divergences. Due to the technical nature of some of the assumptions necessary on $K_{\sigma}$ for this result, we defer a proper presentation to Section A.2 and present only an informal version of the result below. The proof relies on uniform convergence results for kernel regression applied to the individual terms in (4.1).
Theorem 5.1.
Estimation and plotting details. We estimate the term $\mathbb{E}\left[d_{\phi}(\pi_{Y\mid X},g(X))\mid h(X)\right]$ in (5.1) by replacing the $(x_{i},y_{i})$ data point pairs with $(x_{i},d_{\phi}(e_{y_{i}},g(x_{i})))$ (i.e. the divergence of $g(x_{i})$ from the one-hot encoded label $y_{i}$) and then using (5.2). The same is done with $\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)]$, except the label is replaced with pointwise accuracy. We also estimate the density $p_{h(X)}$ using a kernel density estimate (denominator of (5.2)) with the same kernel, and refer to this estimate as $\hat{p}_{h(X)}$.
We then plot the band $\mathbb{E}[\mathbbm{1}_{Y=c(x)}\mid h(x)]\pm\hat{p}_{h(X)}(h(x))\rho(h(x))/2$ at uniformly spaced values of $h(x)$, and also overlay the plot of $\hat{p}_{h(X)}(h(x))$ after normalizing so that the maximum kernel density value is 1. Scaling (5.1) by $\hat{p}_{h(X)}(h(x))$ is important for making it visually obvious where most of the sharpness error is concentrated, and the overlayed density plot allows the same for the calibration error. We point out that another common pitfall in several prior works is to omit the confidence histograms or density plots when using reliability diagrams to visualize calibration, and this makes it not possible to accurately compare the calibration of two models.
Experiments. To illustrate our method, we revisit the experiments of Section 4.2. In Figure 1 we display the calibration-sharpness diagrams for the baseline ViT model as well as each recalibration method, under the exact same settings as Section 4.2, for the choice of MSE/Brier score since some of the recalibration approaches have unbounded KL. Our kernel regression estimates were computed using a Gaussian kernel with $\sigma=0.05$; we discuss different choices of kernel and bandwidth in Appendix B. Due to the size of the dataset, we also replace the full kernel regression estimate of (5.2) with a mean over 10 estimates in which we randomly subsample the data to 5000 data points, so as to fit the entire computation into memory (we use a single A5000 GPU). The standard deviations of these estimates is shown alongside the mean calibration error scaled by 100 ($d_{\varphi,\mathrm{CAL}}$ in the plot); in all cases they are negligible, and we checked that subsampling loses virtually nothing. We also display the estimate of the full divergence ($d_{\phi,\mathrm{TOT}}$), which is computed without subsampling. Lastly, for stylistic reasons, we opt to threshold the curve defining the bottom of the sharpness band at 0, but do not threshold the top curve at 1.
(a) ViT (Baseline)
(b) ViT + TS
(c) ViT + HB
(d) ViT + IR
(e) ViT + MRR
Figure 1: Proposed calibration-sharpness diagrams using MSE for the experiments of Section 4.2.
6 Conclusion
Limitations. Our work has advocated for the reporting of proper scoring rules and calibration metrics implied by such rules. However, we have not addressed the question of which rules are most appropriate for which settings, i.e. MSE/Brier score vs. KL divergence. An investigation into the trade-offs between these and similar choices would be an interesting avenue for future work. Furthermore, we have provided only one implementation of our viewpoint in Section 5, where we estimate calibration and sharpness error using kernel regression and visualize both in a reliability diagram format. We believe constructing good visualizations for calibration and related quantities is still an area that could be developed further. Lastly, while we prove a consistent estimation result in the form of Proposition 5.1, there is still much work to be done on analyzing optimal convergence rates for estimators of calibration error.
Broader Impacts. While calibration has become an increasingly important topic given the use of machine learning models for decision-making, our work concerns only the theory and practice of reporting calibration. As a result, we do not anticipate any negative societal impacts as immediate consequences of this work.
Acknowledgements
Rong Ge and Muthu Chidambaram are supported by NSF Award DMS-2031849 and CCF-1845171 (CAREER).
Appendix A Omitted Results and Proofs
Here we include all technical results and proofs omitted from the main text.
A.1 Proofs for Section 4.2
@@ Appendix A Omitted Results and Proofs
$\mathbb{E}[d_{\phi}(\mathbb{E}[\pi_{Y\mid X}\mid g(X)],g(X))]\geq\mathbb{E}[d_{\varphi}(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid g(X)],h(X))].$ (A.2)
@@
$\mathbb{E}[d_{\varphi}(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid g(X)],h(X))]$ $=\mathbb{E}[\varphi(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid g(X)])]-\mathbb{E}[\varphi(h(X))]$
$\quad+\mathbb{E}[\varphi^{\prime}(h(X))(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid g(X)]-h(X))]$
$\geq\mathbb{E}[\varphi(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)])]-\mathbb{E}[\varphi(h(X))]$
$\quad+\mathbb{E}[\varphi^{\prime}(h(X))(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)]-h(X))]$ (A.3)
@@
The proof follows identical steps to that of Lemma 4.5, except that we apply the data processing inequality to obtain (A.2), which is justified since $c(X)$ is fixed after conditioning on $g(X)$. ∎
A.2 Consistency of Calibration Error Estimation
@@
Assumption A.1 (Kernel Assumptions).
@@
Assumption A.2 (Distributional Assumptions).
@@
Theorem A.3 (Theorem 2 in Devroye (1978), Paraphrased).
Let $K$ be a kernel satisfying Assumption A.1 with a sequence of bandwidths $\sigma_{n}$, and suppose $\pi$ satisfies Assumption A.2. Then if $\sigma_{n}\to 0$ and $n\sigma_{n}^{2d}/\log n\to\infty$, we have uniform convergence of the kernel regression estimator $\hat{m}_{\sigma_{n}}$:
@@
$\hat{d}_{\varphi,n}=\frac{1}{n}\sum_{i=1}^{n}\varphi\big{(}\hat{m}_{\sigma_{n}}(h(x_{i}))\big{)}-\varphi\big{(}h(x_{i})\big{)}-\varphi^{\prime}\big{(}h(x_{i})\big{)}\big{(}\mathbbm{1}_{y_{i}=c(x_{i})}-h(x_{i})\big{)}.$ (A.5)
Suppose $K$ satisfies Assumption A.1 and $\sigma_{n}$ satisfies $\sigma_{n}\to 0$ and $n\sigma_{n}^{2}/\log n\to\infty$. If we additionally have that $E[\mathbbm{1}_{Y=c(X)}\mid h(X)=h(x)]$ is continuous in $h(x)$ and that $h(X)$ has a density bounded away from zero, it follows that:
$\hat{d}_{\varphi,n}\to d_{\varphi}\left(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)],h(X)\right)$ (A.6)
@@
$\quad-\mathbb{E}[\varphi^{\prime}(h(X))(\mathbbm{1}_{Y=c(X)}-h(X))].$ (A.7)
It is immediate from the law of large numbers that the second and third terms in (A.5) converge to their counterparts in (A.7). It remains to show the convergence of the first term of (A.5).
We observe that since $\mathbbm{1}_{Y=c(X)}\leq 1$, along with the fact that we have assumed $\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)=h(x)]$ is continuous in $h(x)$ and that $h(X)$ has a density bounded away from zero, we satisfy all of the conditions of Assumption A.2. Thus we have uniform convergence of $\hat{m}_{\sigma_{n}}$ from Theorem A.3.

### v2

@@ Appendix A Background Results
$\mathbb{E}[d_{\phi}(\mathbb{E}[Z\mid X],X)]$ $=\mathbb{E}[\phi(\mathbb{E}[Z\mid X])]-\mathbb{E}[\phi(X)]-\mathbb{E}[\nabla\phi(X)^{\top}(\mathbb{E}[Z\mid X]-X)]$
$=\mathbb{E}[\phi(\mathbb{E}[Z\mid X])]-\mathbb{E}[\phi(X)]-\mathbb{E}[\nabla\phi(X)^{\top}(Z-X)].$ (A.1)
@@
See 4.2
@@
We first note that the case of $q=1$ is not a Bregman divergence since Bregman divergences are differentiable in their first argument. For $q>1$, Lemma 2 in Boissonnat et al. (2010) shows that a twice differentiable function $\phi$ induces a symmetric Bregman divergence $d_{\phi}$ iff the Hessian of $\phi$ is constant. This implies that for any $L^{q}$ norm (raised to the $q$ power) to be a Bregman divergence, differentiating it twice in the first argument must yield a constant. It is easy to see that this is only the case for $q=2$, and the result follows. ∎
Appendix B Proofs for Section 4.1
See 4.5
Proof.
@@
$d_{\phi}(\mathbb{E}[\pi_{Y\mid X}\mid g(X)],g(X))$ $\geq d_{\varphi}(\mathbb{E}[\pi_{Y\mid X}\mid g(X)]_{c(X)},g(X)_{c(X)})$
$=d_{\varphi}(e_{c(X)}^{\top}\mathbb{E}[\pi_{Y\mid X}\mid g(X)],e_{c(X)}^{\top}g(X))$
$=d_{\varphi}(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid g(X)],h(X)).$ (B.2)
@@
$\mathbb{E}[d_{\varphi}(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid g(X)],h(X))\mid h(X)]$ $=\mathbb{E}[\varphi(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid g(X)])\mid h(X)]-\varphi(h(X))$
$\quad-\mathbb{E}[\varphi^{\prime}(h(X))(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid g(X)]-h(X))\mid h(X)]$
$\geq\varphi(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)])-\varphi(h(X))$
$\quad-\varphi^{\prime}(h(X))(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)]-h(X))$ (B.3)
@@
We need only show that $d_{\phi}(x,y)\geq d_{\varphi}(x_{i},y_{i})$ for the choices of KL divergence and logistic loss. Let us as usual identify $x$ and $y$ with vectors in $\mathbb{R}^{k}$ corresponding to distributions in $\Delta^{k-1}$. Then, fixing an $i\in[k]$, the map $f(x)=\mathrm{Bernoulli}(x_{i})$ is a Markov kernel, and by the data-processing inequality we have $d_{\phi}(x,y)\geq d_{\phi}(f(x),f(y))$. Since $d_{\phi}(f(x),f(y))=d_{\varphi}(x_{i},y_{i})$, the result follows. ∎
Appendix C Proofs for Section 4.2
See 4.8
Proof.
From Lemma 4.1, we have:
$\mathbb{E}\left[d_{\phi}(\pi_{Y\mid X},g(X))\mid h(X)=p\right]\geq\mathbb{E}\left[d_{\phi}(\mathbb{E}[\pi_{Y\mid X}\mid g(X)],g(X))\mid h(X)=p\right].$ (C.1)
Applying Lemma 4.6 then yields the result. ∎
@@
Assumption C.1 (Kernel Assumptions).
@@
Assumption C.2 (Distributional Assumptions).
@@
Theorem C.3 (Theorem 2 in Devroye (1978), Paraphrased).
Let $K$ be a kernel satisfying Assumption C.1 with a sequence of bandwidths $\sigma_{n}$, and suppose $\pi$ satisfies Assumption C.2. Then if $\sigma_{n}\to 0$ and $n\sigma_{n}^{2d}/\log n\to\infty$, we have uniform convergence of the kernel regression estimator $\hat{m}_{\sigma_{n}}$:
@@
$\hat{d}_{\varphi,n}=\frac{1}{n}\sum_{i=1}^{n}\varphi\big{(}\hat{m}_{\sigma_{n}}(h(x_{i}))\big{)}-\varphi\big{(}h(x_{i})\big{)}-\varphi^{\prime}\big{(}h(x_{i})\big{)}\big{(}\mathbbm{1}_{y_{i}=c(x_{i})}-h(x_{i})\big{)}.$ (C.3)
Suppose $K$ satisfies Assumption C.1 and $\sigma_{n}$ satisfies $\sigma_{n}\to 0$ and $n\sigma_{n}^{2}/\log n\to\infty$. If we additionally have that $E[\mathbbm{1}_{Y=c(X)}\mid h(X)=h(x)]$ is continuous in $h(x)$ and that $h(X)$ has a density bounded away from zero, it follows that:
$\hat{d}_{\varphi,n}\to\mathbb{E}[d_{\varphi}\left(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)],h(X)\right)]$ (C.4)
@@
$\quad-\mathbb{E}[\varphi^{\prime}(h(X))(\mathbbm{1}_{Y=c(X)}-h(X))].$ (C.5)
It is immediate from the law of large numbers that the second and third terms in (C.3) converge to their counterparts in (C.5). It remains to show the convergence of the first term of (C.3).
We observe that since $\mathbbm{1}_{Y=c(X)}\leq 1$, along with the fact that we have assumed $\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)=h(x)]$ is continuous in $h(x)$ and that $h(X)$ has a density bounded away from zero, we satisfy all of the conditions of Assumption C.2. Thus we have uniform convergence of $\hat{m}_{\sigma_{n}}$ from Theorem C.3.